A Complete Bibliography of Electronic Publishing—Origin, Dissemination, and Design

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

11 May 2013 Version 2.22

This bibliography summarizes the journal contents. A UNIX refer version and related material, including style files for TEX, L^{A}TEX, BibTEX, and nroff/troff, and sample articles, can be found on the machine ftp.cs.

nott.ac.uk in the directory ep/epodd.

The journal is also available on CD ROM in Adobe Acrobat format (a high-quality document viewing system with hypertext links and indexing); see the CAJUN Project description [SBE^{+}93].

Title word cross-reference

4 [BH91].

8879 [Bar92].

A. [Wei92]. accuracy [Sen93]. Acrobat [SBE^{+}93]. Active [TB90, ET94, QV94]. activity [JS90]. adaptation [BB89a]. adaptive [ZB93]. Adobe [Har93]. advanced [ABH93]. aid [PD90]. algorithm [IO93]. ALIVE [PCS92]. alternative [Har93]. Amsterdam [WV89]. Analysing [HH93]. Analysis [LB95, BAC93]. André [Wei92]. angles [Knu90]. Antoni [Pic94]. appearance [Bro91]. Apple [BH91]. Applications [Bro95, vES95, Kac95]. approach [FX93, HI93, Kac95, SD91]. Arabic [Har95, SB92]. architecture [HGC94]. Author [Ano95b, Ano95a]. Authoring [BB88a, Fei90]. auto [HH94]. auto-hinting [HH94]. Automatic [Aho95, Kok94, SBA92, PM93]. Automatically [FPS89]. aware [Bro95]. awareness [Man93]. backed [BO89]. Based [SB95, ZD95, AHW89, CK95, FW93, Han90, Kok94, MBM93, Ric94]. Bases [Luc88]. basic [Fur92]. Bayesian [SD91]. be [JS90]. between [LEB93, HS93]. bi [HB90]. bi-directional [HB90]. bibliography [HM89]. Block [SB95]. Block-Based [SB95]. Book [Bar92, Bar94, Pic92, Pic94, Rah92b, Rah92a, Wei92]. Books [Bar94]. both [Mur95]. Braille [AJ88]. browsing [PD90].

C [GP89]. C-like [GP89]. CAJUN [SBE^{+}93]. Call [Zap94]. Can [And89, JS90]. Canto [NR93]. CD [Bar94]. CD-ROM [Bar94]. character [HH93, IO93]. characters [AV94, PM93]. Chinese [PM93]. Collier [Pic92]. Colophon [AS93, BK95, HM93]. communication [BJ92]. compilation [GB93]. components [HM91]. Compression [ZD95]. computational [BCD89]. Computer [BB88c, BU92]. Computerized [AJ88]. concepts [ABH93]. concerns [LL93]. Conference [Wei92, AGS93, BBKF95, QHM93]. Considerations [MB91]. constraint [MBM93]. constraint-based [MBM93]. constraints [Erf93, Sen93]. containing [Mur95]. Content [Aho95]. Contents [Ano95e]. Context [Bro95]. Context-aware [Bro95]. control [Bro91, SB93]. Convergent [Dod90]. Cookbook [Rei88]. Coordinate [Dür93]. Coordinate-independent [Dür93]. coursework [BJ92]. crashes [And89]. Creating [Bro95]. creation [RMF93]. Crookes [Rah92b]. Curriculum [Ham92, Dys92]. curve [IO93, Kok94]. cutting [Kar91].

data [ACG91, NR93]. Database [RSH90, ABH93]. Databases [SF91]. David [Pic92]. Definitions [MB91]. Description [ON88, Dür93]. Design [Pic92, SLH^{+}93, BH93]. Desktop [Bar94, Pic92]. Developing [Man93]. Development [ON88]. device [BB89a]. device-independent [BB89a]. differences [LEB93]. diffusion [ZB93]. Digital [AGS93, DM89, Kar91, SB93, AH92]. digitized [Knu90]. Diller [Pic94]. directed [CK95]. directional [BB89a, HB90]. discrimination [Dys93]. display [Dys93]. dissemination [Har94b]. distributed [HGC94, PCS92, HGC94]. ditroff [SB92]. ditroff/ffortid [SB92]. Do [Bro89]. Document [BE89, Bro88a, Bro95, BBKF95, CK95, Lev93, LB95, Luc88, MB91, Pic94, QHM93, ZD95, Arn93, FX93, Fur92, HM91, HI93, Rad90]. documentation [PCS92]. Documents [Bro89, Fei90, FQA88, Hee92, Kin95, KS95, Mun95, RA95, SF91, WvV91, AHW89, BM91, Bar95, BCD89, CB92, CMPdVS91, ET94, Erf93, FW93, FL93, FPS89, GB93, HS93, KP93, KP95, Mur95, QV94, RV93, TB90, HGC94]. Domains [KFA95]. Drake [Zap94]. Drawing [BKW89]. Dynamic [AV94, Sta93a].

Editing [CB92, FQA88]. Editor [WW95, CMPdVS91, HB90, MBM93, Ric94]. Editorial [And92, AGS93, BB88b, BB88d, BB88e, BB89b, BB89c, BB89d, BB89e, BF90a, BF90b, BF90c, BF90d, BF91a, BF91b, BF91c, BF92a, BF92b, BF93a, BF93b, BF95, BBKF95, FB94, Hal94, Her94, Luc91, QHM93, Qui94]. editors [Wei92]. efficient [HI93]. Electronic [Bar94, BB88c, BBKF95, HHR94, Ham92, QHM93, Rei88, BCD89, BJ92, BU92, McC92, Ric94, RMF93]. Embedded [BB95]. encoded [BM91]. ends [CB92, Ham92]. enhancements [CCT94]. Environment [KS95, Won94]. EP [BB88a, BJ92, CB92, Ham92]. EP-odd [BB88a]. EP-odds [CB92, Ham92]. EP’94 [QHM93]. EP’96 [BBKF95, Ano95b, Ano95c]. Error [ZB93]. European [Wei92]. Examining [JS90]. example [Dür93, HHR94, McC92]. existing [Kac95]. Experiment [BE89, Rei88]. exploration [TB90]. Extended [Har95]. extensibility [Sta93b]. extraction [PM93].

Facet [All95]. ffortid [SB92]. File [Mur95]. Files [LB95, CEMW91, SBA92]. Filtering [KS95]. First [Wei92]. fitting [IO93]. font [BH93, Dür93, Kla93, Kok94, MB93, Sta93b, ZI93]. font-scaler [Sta93b]. Fonts [BH91, DM95, Har93, HP93, IO93, Sta93a, Har93]. format [Mur95]. Formatter [WW95]. formatters [And89]. Formatting [BE89, Han90, BB89a, SB92]. Forum [SGBO^{+}93]. foundation [Zap94]. Framework [Bro95]. Francis [Zap94]. full [AHW89, HB90]. full-screen [HB90]. full-text [AHW89]. Function [Han90]. Function-based [Han90]. functional [FS90].

Gabriel [Bar94]. Generation [Aho95, Kok94, MB93]. Glyphs [Har95]. good [JS90]. grammar [FW93, KP93]. grammar-based [FW93]. Granularity [Hee92]. Grif [Ric94]. Guide [Bar94, Bar92].

Halftoning [SB93]. handwriting [DM95]. hard [BO89]. hard-backed [BO89]. HB1 [SLH^{+}93]. Helvetica [LEB93]. hinting [HH94, PM93]. history [Fur92]. Hybrid [Dod90]. hyper [JS90]. hyper-activity [JS90]. hyperbase [SLH^{+}93]. Hypermedia [Fei90, Kac95, RML94]. Hypertext [Bro88b, Bro89, CCT94, Pol90, Rad90, SF91, Bro91, BB95, Dys92, FS90, HGC94, HS93, JS90, NR93, PD90, Ric94, SD91, FPS89, Wei92]. hypertexts [RSH90]. HyTime [CBDH94, Erf93]. hz} [Zap93]. hz}-program [Zap93].

IGD [Fei90]. Images [ZD95, HI93].

Imaging [AGS93]. Implementation [ON88, SLH^{+}93, WV89]. Implications [LB95]. Importance [DM89]. Important [Fur92]. Improving [Dys93, Sen93, Kae90]. incremental [FX93]. independent [BB89a, Dür93]. Index [Ano95b, Ano95c, Ano95a, Ano95d, Ric94]. Indexes [BK88]. indexing [LL93]. Individualized [HHR94]. industrial [Won94]. Infinifont [MB93]. Information [SD91, ACG91]. InScript [GP89]. integrated [HM89]. intelligent [Sta93a]. Interactive [KFA95, SGBO^{+}93]. Interactively [FQA88]. interchangeability [NW92]. interface [CMPdVS91, HS93]. Interfaces [KFA95]. Interleaf [ET94]. Introduction [Rah92a]. ISO [Bar92]. Issue [BBKF95, AGS93, BB95, QHM93]. Issues [ACG91].

J [Wei92]. Joan [Bar92]. Journal

[Dod90, SBE^{+}93, SGBO^{+}93]. Journals [Bar94].

Kanji [Dür93]. Key [Ano95c]. Keyword [Ano95d]. Khmer [Har94a]. Knowledge [KFA95]. Knowledge-rich [KFA95].

Language [Mun95, Arn93]. Languages [ON88]. Large [Fei90, Luc88, RSH90]. Layout [Bro88a, Mur95]. letterforms [Man93]. level [BAC93]. ligatures [Kok94]. like [GP89]. Line [Pic94]. linguistic [MBM93]. link [PCS92]. Linking [Bro88b]. Literature [Bar94]. live [PCS92]. live-link [PCS92]. logical [BAC93, Bro91, HI93, Mur95, RV93]. Lossy [ZD95]. low [BAC93, Dys93]. low-level [BAC93].

M [Bar92]. Mail [Bar94]. Making [QV94]. management [SLH^{+}93]. manipulating [CMPdVS91]. Manipulation [BBKF95, QHM93, FL93]. Maps [Bro89]. Mark [AJ88, BB95]. Mark-up [AJ88, BB95]. Marking [BJ92]. MarkItUp [FX93]. Master [Har93]. matching [HH93]. Mediating [HS93]. Merging [RV93]. meta [FS90]. meta-structure [FS90]. Methods [LB95]. MHEG [RML94]. Michael [Bar94]. micro [Zap93]. micro-typography [Zap93]. mixed [HI93]. Model [Han90, BM91, BCD89, NR93]. Modelling [Kin95, ACG91]. Models [Aho95, FS90]. Multicard [RML94]. Multimedia [Kin95, CCT94, Erf93, Har94b, SGBO^{+}93]. Multiple [Har93]. music [FL93].

N [Wei92]. Navigate [Bro89]. Nearest [CEMW91, AHW89]. Nearest-neighbour [CEMW91]. Need [Bro89]. neighbor [AHW89]. neighbour [CEMW91]. net [Rad90]. network [RMF93]. networks [CEMW91, SD91]. news [SvR91]. Newspaper [HHR94]. Nicely [BKW89]. Node [PD90]. non [Kac95]. non-hypermedia [Kac95]. Norbert [Rah92a]. note [Knu90]. Notes [BH91]. NRT [SvR91]. Numbering [HM91].

Object [FFE95, Sta93b, RMF93]. Object-orientation [Sta93b]. objects [Bro91]. ODA [NW92]. odd [BB88a]. odds [CB92, Ham92]. online [Pol90, SGBO^{+}93]. open [HGC94]. Optical [ZI93, AV94]. optimization [Kok94]. orientation [Sta93b]. oriented [JS90]. outline [PM93, Sta93a]. outlines [HP93].

Page [ON88]. Pagination [BKKW95]. Paper [Dod90]. papers [Fur92]. Paradigms [FFE95, TB90]. Paragraph [AHW89]. Paragraph-based [AHW89]. Parallel [BE89, Bro88a]. parametric [MB93]. Parametrization [Har93]. Parser [WV89]. partial [Bar95]. Pattern [ZD95]. Pattern-Based [ZD95]. PDF [LB95, SB95].

HyTime [FFE95]. Performance [LEB93]. Phase [DM89]. Philip [Rah92b]. physical [RV93]. PIC [BE89]. plus [Dod90]. popularity [PD90]. Portable [Bar95]. possible [Kla93]. PostScript [GP89, Har93]. preparation [Fur92, MB91]. preprocessor [GP89]. Presentation [Mun95, Pol90]. prevent [And89]. Printing [BK88]. problems [Bar95, CB92]. Proceedings [AGS93, BBKF95, QHM93]. process [Kac95]. Processing [BE89, Bro88a, WvV91, HM89, Pic94]. profiles [ZI93]. program [Zap93]. project [SBE^{+}93]. projection [ZI93]. properties [BCD89]. prospects [ABH93]. providing [Kac95]. Publication [Dod90, HHR94]. Publisher [Rah92b]. Publishing [Bar94, BB88c, BBKF95, Ham92, QHM93, Rei88, ABH93, BU92, McC92, SBE^{+}93]. punch [Kar91].

Qualitative [BAC93]. queries [Arn93].

R [Bar94]. Random [DM95]. Raster [AGS93, ZB93]. Rasterizing [HP93]. reading [LEB93]. recognition [FX93, HI93, Sen93, ZI93]. Reconsidered [BKKW95]. reference [RMF93]. regularisation [Sta93a]. regularly [FPS89]. Repositories [FFE95]. Representation [FL93, Kla93]. representing [RMF93]. resolution [Dys93]. Results [LB95]. Retrieval [All95, ACG91, SvR91, SD91]. reuse [Lev93, Rad90]. Review [Bar92, Bar94, Pic92, Pic94, Rah92b, Rah92a, Wei92, Har94b]. rich [KFA95]. Rita [CMPdVS91]. Rizk [Wei92]. Robert [Bar92]. role [Rad90]. ROM [Bar94]. Round [Bro89]. Rules [Pic92].

scaler [Sta93b]. scaling [AV94]. Scheme [ZD95]. scholars [MBM93]. Schwartz [Rah92a]. Science [BB88c]. scientists [BU92]. Scottish [McC92]. screen [BO89, HB90]. Scrimshaw [Arn93]. script [Kok94]. Search [Luc88]. Searching [Bro88b, AHW89, CEMW91, Pol90]. semantic [Rad90]. Separate [GB93, BB95]. Separation [LL93]. services [Kac95]. Set [Har95]. SGML [Aho95, Bar89, Bar92, FFE95, Kae90, MB91, NW92, WV89, WvV91, vES95]. SGML/ HyTime [FFE95]. shapes [HH93, HH94]. Sheets [RA95]. signatures [CEMW91]. SIMON [FW93]. simulation [DM95]. Sir [Zap94]. sized [AV94]. Smith [Bar92]. smoothness [Kok94]. Society [Zap94]. solutions [Bar95, CB92]. Some [BCD89]. sources [Fur92]. Spaces [All95]. Special [AGS93, BBKF95, QHM93]. Specification [Erf93]. Spectra [DM89]. splines [Kla93]. standard [RML94]. Standardization [ON88]. Standards [AJ88]. Stick [Bro95]. Stick-e [Bro95]. Strategy [Luc88]. Streitz [Wei92]. string [HH93]. Structure [PM93, FX93, FS90, HI93, RMF93]. Structured [FQA88, Hee92, KS95, Mun95, RA95, SB95, And89, CB92, CMPdVS91, FW93, FPS89, GB93, HS93, KP93, KP95, QV94, Ric94]. structures [BAC93, Mur95, RV93]. structuring [SBA92]. Stutely [Bar92]. support [RSH90]. symbols [Dys93]. SYNDOC [KS95]. syntax [CK95]. syntax-directed [CK95]. system [CCT94, FW93, HH94, MB93, PCS92, RML94, SLH^{+}93]. systems [BB88a, FS90, Fur92, Har94b, Lev93, SD91, Sen93].

Tabular [WW95]. tagging [BM91]. task [LEB93]. Teaching [AH92, BU92, McC92]. Techniques [Pic94, HH93]. technology [Har94b]. Telecommunications [Dod90]. temporal [Erf93]. text [AHW89, BB89a, CEMW91, SBA92, Sen93]. texture [SB93]. thesauri [Pol90]. Times [LEB93]. Tioga [TB90]. Tips [Pic94]. tool [SvR91]. Tools [BK88]. toolset [BM91].

Traditional [Har95]. train [And89].

training [JS90]. training-oriented [JS90]. Transformation [KP93, KP95, CK95, FW93]. transformations [Arn93]. transforming [FPS89]. translation [CK95]. transputer [CEMW91]. tree [CK95]. Trees [BKW89]. tri [BB89a]. tri-directional [BB89a]. triroff [BB89a]. troff [BB89a]. two [TB90]. Type [DM89, MB91]. Typecase [Har95]. Typesetting [AJ88, Har94a]. Typographic [RA95, HH94]. typographical [Sen93]. typographics [Man93]. Typography [AGS93, BBKF95, QHM93, AH92, Zap93, Pic92].

Unicode [BH93, Har95]. uniform [BM91]. universal [HH94]. usability [JS90]. Use [Bar89, CBDH94, KP93]. used [Pol90]. USENET [Rei88]. user [Bar92, CMPdVS91]. Using [BE89, Bro91, RML94, ABH93, BJ92, CEMW91, Dür93, Erf93, SD91, Sen93].

Variable [Kla93, AV94]. variable-sized [AV94]. Ventura [Rah92b]. version [HB90]. very [RSH90]. vi [HB90]. Vi.iv [HB90]. View [AJ88]. Volume [Ano95e]. VORTEXT [BO89].

Weaving [RMF93]. web [RMF93, vES95]. width [Kla93]. Within [Bro88b]. Word [Ano95c]. work [RMF93]. writing [Rad90].

Xtable [WW95].

References

Aberer:EPODD-6-4-469

[ABH93] Karl Aberer, Klemens Böhm, and Christoph Hüser. The prospects of publishing using advanced database concepts.

Electronic Publishing—Origination, Dissemination, and Design, 6(4):469–480, December 1993. CODEN EPODEU. ISSN 0894-3982.

Agosti:EPODD-4-4-219

[ACG91] M. Agosti, R. Colotti, and G. Gradenigo. Issues of data modelling in information retrieval. Electronic Publishing—Origination, Dissemination, and Design, 4(4):219– 237, December 1991. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-6-3-115

[AGS93] Jacques André, Jakob Gonczarowski, and Richard Southall. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference. Electronic Publishing—Origination, Dissemination, and Design, 6(3):115– 116, September 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-5-2-79

[AH92] Jacques André and Roger D.

Hersch. Teaching digital typography. Electronic Publishing —Origination, Dissemination, and Design, 5(2):79–89, June 1992. CODEN EPODEU. ISSN 0894-3982.

Ahonen:EPODD-8-2/3-195

[Aho95] Helena Ahonen. Automatic generation of SGML content models. Electronic Publishing —Origination, Dissemination, and Design, 8(2/3):195–206, June/September 1995. CODEN EPODEU. ISSN 0894-3982. Al-Hawamdeh:EPODD-2-4-179

[AHW89] Suliman Al-Hawamdeh and Peter Willett. Paragraph-based nearest neighbor searching in full-text documents. Electronic Publishing—Origination, Dissemination, and Design, 2(4): 179–192, December 1989. CO-DEN EPODEU. ISSN 0894-3982. Arrabito:EPODD-1-2-117

[AJ88] R. Arrabito and H. Jürgensen. Computerized Braille typesetting: Another view of markup standards. Electronic Publishing—Origination, Dissemination, and Design, 1(2):117– 132, September 1988. CODEN EPODEU. ISSN 0894-3982. Allen:EPODD-8-2/3-247

[All95] R. B. Allen. Retrieval from facet spaces. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 247–257, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-2-3-169

[And89] Jacques André. Can structured formatters prevent train crashes? Electronic Publishing —Origination, Dissemination, and Design, 2(3):169–173, October 1989. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-5-2-53

[And92] Jacques André. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 5 (2):53–54, June 1992. CODEN EPODEU. ISSN 0894-3982.

Anonymous:EPODD-8-4-369

[Ano95a] Anonymous. Author index.

Electronic Publishing—Origination, Dissemination, and Design, 8(4):369, December 1995. CODEN EPODEU. ISSN 0894-3982.

Anonymous:EPODD-8-2/3-275

[Ano95b] Anonymous. EP’96 author index. Electronic Publishing —Origination, Dissemination, and Design, 8(2/3):275, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Anonymous:EPODD-8-2/3-277

[Ano95c] Anonymous. EP’96 key word index. Electronic Publishing —Origination, Dissemination, and Design, 8(2/3):277, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Anonymous:EPODD-8-4-371

[Ano95d] Anonymous. Keyword index.

Electronic Publishing—Origination, Dissemination, and Design, 8(4):371–372, December 1995. CODEN EPODEU. ISSN 0894-3982.

Anonymous:EPODD-8-4-i

[Ano95e]  Anonymous.  Volume con-

tents. Electronic Publishing —Origination, Dissemination, and Design, 8(4):i–iii, December 1995. CODEN EPODEU. ISSN 0894-3982.

Arnon:EPODD-6-4-385

[Arn93] Dennis S. Arnon. Scrimshaw: A language for document queries and transformations. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):385–396, December 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-6-3-307

[AS93] Jacques André and Richard Southall. Colophon. Electronic Publishing—Origination, Dissemination, and Design, 6 (3):307–308, September 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:EPODD-7-4-231

[AV94] J. André and I. Vatton. Dynamic optical scaling and variable-sized characters. Electronic Publishing—Origination, Dissemination, and Design, 7 (4):231–250, December 1994. CODEN EPODEU. ISSN 0894-3982.

Belaid:EPODD-6-4-435

[BAC93] Abdel Belaïd, Julian C. Anigbogu, and Yannich Chenevoy. Qualitative analysis of low-level logical structures. Electronic Publishing—Origination, Dissemination, and Design, 6(4): 435–446, December 1993. CO-DEN EPODEU. ISSN 0894-3982.

Barron:EPODD-2-1-3

[Bar89]  D. W. Barron.  Why use

SGML? Electronic Publishing —Origination, Dissemination, and Design, 2(1):3–24, April 1989. CODEN EPODEU. ISSN 0894-3982.

Barron:EPODD-5-3-157

[Bar92] David W. Barron. Book review:

Joan M. Smith and Robert Stutely, SGML: the user’s guide to ISO 8879. Electronic Publishing—Origination, Dissemination, and Design, 5(3):157, September 1992. CODEN EPODEU. ISSN 0894-3982.

Barron:EPODD-7-1-49

[Bar94] David W. Barron. Book review:

Michael R. Gabriel, A Guide to the Literature of Electronic Publishing: CD-ROM, Desktop Publishing, and Electronic Mail, Books and Journals. Electronic Publishing—Origination, Dissemination, and Design, 7 (1):49, March 1994. CODEN EPODEU. ISSN 0894-3982.

Barron:EPODD-8-4-343

[Bar95] David W. Barron. Portable documents: problems and (partial) solutions. Electronic Publishing—Origination, Dissemination, and Design, 8(4):343– 367, December 1995. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-0-0-4

[BB88a] D. F. Brailsford and R. J.

Beach. Authoring systems for EP-odd. Electronic Publishing —Origination, Dissemination, and Design, 0(0):4–9, January 1988. CODEN EPODEU. ISSN 0894-3982. This is a pilot issue of the journal.

Brailsford:EPODD-0-0-1

[BB88b] D. F. Brailsford and R. J. Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 0 (0):1–3, January 1988. CO-DEN EPODEU. ISSN 0894-3982. This is a pilot issue of the journal.

Brailsford:EPODD-0-0-13

[BB88c] D. F. Brailsford and R. J. Beach. Electronic publishing and computer science. Electronic Publishing—Origination, Dissemination, and Design, 0 (0):13–21, January 1988. CO-DEN EPODEU. ISSN 0894-3982. This is a pilot issue of the journal.

Brailsford:EPODD-1-1-1

[BB88d] David F. Brailsford and Richard J. Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 1(1): 1–2, April 1988. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-1-2-77

[BB88e] David F. Brailsford and Richard J. Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 1(2): 77–78, September 1988. CO-DEN EPODEU. ISSN 0894-3982.

Becker:EPODD-2-3-119

[BB89a] Zeev Becker and Daniel Berry. triroff, an adaptation of the device-independent troff for formatting tri-directional text. Electronic Publishing—Origination, Dissemination, and Design, 2(3):119–142, October 1989. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-2-1-1

[BB89b] David F. Brailsford and Richard J.

Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 2(1): 1–2, April 1989. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-2-2-63

[BB89c] David F. Brailsford and Richard J.

Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 2(2): 63–64, July 1989. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-2-3-117

[BB89d] David F. Brailsford and Richard J.

Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 2(3): 117–118, October 1989. CO-DEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-2-4-177

[BB89e] David F. Brailsford and Richard J.

Beach. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 2(4): 177–178, December 1989. CO-DEN EPODEU. ISSN 0894-3982.

Brown:EPODD-8-1-1

[BB95] P. J. Brown and Heather Brown. Embedded or separate hypertext mark-up: is it an issue? Electronic Publishing —Origination, Dissemination, and Design, 8(1):1–13, March 1995. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-8-2/3-49

[BBKF95] Allen Brown, Anne Brüggemann-Klein, and An Feng. Editorial: Special issue: Proceedings of EP’96, the Electronic Publishing, Document Manipulation, and Typography Conference. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):49–50, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Bench-Capon:EPODD-2-4-231

[BCD89] Trevor J. M. Bench-Capon and Paul E. Dunne. Some computational properties of a model for electronic documents. Electronic Publishing—Origination, Dissemination, and Design, 2 (4):231–256, December 1989. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-2-1-25

[BE89] D. F. Brailsford and D. R. Evans. Parallel processing in document formatting: An experiment using PIC. Electronic Publishing—Origination, Dissemination, and Design, 2 (1):25–46, April 1989. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-3-1-1

[BF90a] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 3(1): 1–2, February 1990. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-3-2-63

[BF90b] David F. Brailsford and Richard K.

Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 3(2): 63–64, May 1990. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-3-3-123

[BF90c] David F. Brailsford and Richard K.

Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 3(3): 123–124, August 1990. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-3-4-177

[BF90d] David F. Brailsford and Richard K.

Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 3(4): 177–178, November 1990. CO-DEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-4-1-1

[BF91a] David F. Brailsford and Richard K.

Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 4(1): 1–2, March 1991. CODEN EPODEU. ISSN 0894-3982. Brailsford:EPODD-4-2-61

[BF91b] David F. Brailsford and Richard K.

Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 4(2): 61–62, June 1991. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-4-3-123

[BF91c] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 4(3): 123–124, September 1991. CO-DEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-5-3-103

[BF92a] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 5(3): 103–104, September 1992. CO-DEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-5-4-161

[BF92b] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 5(4): 161–162, December 1992. CO-DEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-6-1-1

[BF93a] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 6(1): 1–2, March 1993. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-6-2-65

[BF93b] David F. Brailsford and Richard K. Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 6(2): 65–66, June 1993. CODEN EPODEU. ISSN 0894-3982.

Brailsford:EPODD-8-4-279

[BF95] David Brailsford and Richard Furuta. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 8 (4):279–280, December 1995. CODEN EPODEU. ISSN 0894-3982.

Bigelow:EPODD-4-3-171

[BH91] Charles A. Bigelow and Kris Holmes. Notes on Apple 4 fonts. Electronic Publishing—Origination, Dissemination, and Design, 4(3):171–181, September 1991. CODEN EPODEU. ISSN 0894-3982.

Bigelow:EPODD-6-3-289

[BH93] Charles Bigelow and Kris Holmes. The design of a Unicode font. Electronic Publishing—Origination, Dissemination, and Design, 6(3):289– 305, September 1993. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-5-2-73

[BJ92] P. J. Brown and R. E. Jones.

Marking EP coursework using electronic communication. Electronic Publishing—Origination, Dissemination, and Design, 5 (2):73–78, June 1992. CODEN EPODEU. ISSN 0894-3982.

Bentley:EPODD-1-1-3

[BK88] J. L. Bentley and B. W.

Kernighan. Tools for printing indexes. Electronic Publishing—Origination, Dissemination, and Design, 1(1):3–18, April 1988. CODEN EPODEU. ISSN 0894-3982.

Bruggemann-Klein:EPODD-8-2/3-273

[BK95] Anne Brüggemann-Klein. Colophon. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):273, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Bruggemann-Klein:EPODD-8-2/3-139

[BKKW95] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. Pagination reconsidered. Electronic Publishing—Origination, Dissemination, and Design, 8 (2/3):139–152, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Bruggemann-Klein:EPODD-2-2-101

[BKW89] A. Brüggeman-Klein and D. Wood. Drawing trees nicely with TEX. Electronic Publishing—Origination, Dissemination, and Design, 2(2):101–115, July 1989. CODEN EPODEU. ISSN 0894-3982.

Barnes:EPODD-4-2-63

[BM91] Julie A. Barnes and Sandra A. Mamrak. A model and toolset for the uniform tagging of encoded documents. Electronic Publishing—Origination, Dissemination, and Design, 4(2): 63–85, June 1991. CODEN EPODEU. ISSN 0894-3982.

Burrill:EPODD-2-3-143

[BO89] Victoria A. Burrill and John A. Ogden. VORTEXT: The hardbacked screen. Electronic Publishing—Origination, Dissemination, and Design, 2(3):143– 156, October 1989. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-1-2-97

[Bro88a] H. Brown. Parallel processing and document layout. Electronic Publishing—Origination, Dissemination, and Design, 1 (2):97–104, September 1988. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-1-1-45

[Bro88b] P. J. Brown. Linking and searching within hypertext. Electronic Publishing—Origination, Dissemination, and Design, 1(1):45–54, April 1988. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-2-2-91

[Bro89] P. J. Brown. Do we need maps to navigate round hypertext documents? Electronic Publishing—Origination, Dissemination, and Design, 2 (2):91–100, July 1989. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-4-2-109

[Bro91] P. J. Brown. Using logical objects to control hypertext appearance. Electronic Publishing —Origination, Dissemination, and Design, 4(2):109–118, June 1991. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-8-2/3-259

[Bro95]  P. J. Brown.  The Stick-

e document: A framework for creating context-aware applications. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 259–272, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Brown:EPODD-5-2-91

[BU92] H. Brown and I. A. Utting. Teaching electronic publishing to computer scientists. Electronic Publishing—Origination, Dissemination, and Design, 5 (2):91–96, June 1992. CODEN EPODEU. ISSN 0894-3982.

Cole:EPODD-5-4-209

[CB92] Fred Cole and Heather Brown. EP-odds and ends: Editing structured documents— problems and solutions. Electronic Publishing—Origination, Dissemination, and Design, 5 (4):209–216, December 1992. CODEN EPODEU. ISSN 0894-3982.

Carr:EPODD-7-3-163

[CBDH94] L. A. Carr, D. W. Barron, H. C. Davis, and W. Hall. Why use HyTime? Electronic Publishing—Origination, Dissemination, and Design, 7(3):163– 178, September 1994. CODEN EPODEU. ISSN 0894-3982.

Clark:EPODD-7-3-147

[CCT94] A. F. Clark, S. L. Cheah, and T. K. Tan. Hypertext and multimedia enhancements to the TEX system. Electronic Publishing—Origination, Dissemination, and Design, 7(3):147– 161, September 1994. CODEN EPODEU. ISSN 0894-3982.

Cringean:EPODD-4-4-185

[CEMW91] Janey K. Cringean, Roger England, Gordon A. Manson, and Peter Willett. Nearestneighbour searching in files of text signatures using transputer networks. Electronic Publishing—Origination, Dissemination, and Design, 4(4):185– 203, December 1991. CODEN EPODEU. ISSN 0894-3982.

Chiba:EPODD-8-1-15

[CK95] Kazuya Chiba and Masaki Kyojima. Document transformation based on syntax-directed tree translation. Electronic Publishing—Origination, Dissemination, and Design, 8(1):15– 29, March 1995. CODEN EPODEU. ISSN 0894-3982.

Cowan:EPODD-4-3-125

[CMPdVS91] D. D. Cowan, E. W. Mackie, G. M. Pianosi, and G. de V. Smit. Rita—an editor and user interface for manipulating structured documents. Electronic Publishing—Origination, Dissemination, and Design, 4(3): 125–150, September 1991. CO-DEN EPODEU. ISSN 0894-3982. See [CB92].

Duan:EPODD-2-1-47

[DM89] G. Duan and R. A. Morris.

The importance of phase in the spectra of digital type. Electronic Publishing—Origination, Dissemination, and Design, 2 (1):47–60, April 1989. CODEN EPODEU. ISSN 0894-3982.

Devroye:EPODD-8-4-281

[DM95] Luc Devroye and Michael Mc-Dougall. Random fonts for the simulation of handwriting. Electronic Publishing—Origination, Dissemination, and Design, 8(4):281–294, December 1995. CODEN EPODEU. ISSN 0894-3982.

Dodd:EPODD-3-1-47

[Dod90] W. P. Dodd. Convergent publication, or the hybrid journal: Paper plus telecommunications. Electronic Publishing —Origination, Dissemination, and Design, 3(1):47–60, February 1990. CODEN EPODEU. ISSN 0894-3982.

Durst:EPODD-6-3-133

[Dür93] Martin J. Dürst. Coordinateindependent font description using Kanji as an example. Electronic Publishing—Origination, Dissemination, and Design, 6(3):133–143, September 1993. CODEN EPODEU. ISSN 0894-3982.

Dyson:EPODD-5-2-63

[Dys92] Mary C. Dyson. The curriculum as hypertext. Electronic Publishing—Origination, Dissemination, and Design, 5(2):63–72, June 1992. CODEN EPODEU. ISSN 0894-3982.

Dyson:EPODD-6-3-231

[Dys93] Mary C. Dyson. Improving discrimination of symbols for display at low resolution. Electronic Publishing—Origination, Dissemination, and Design, 6 (3):231–239, September 1993. CODEN EPODEU. ISSN 0894-3982.

Erfle:EPODD-6-4-397

[Erf93] Robert Erfle. Specification of temporal constraints in multimedia documents using Hy-Time. Electronic Publishing—Origination, Dissemination, and Design, 6(4):397– 411, December 1993. CODEN EPODEU. ISSN 0894-3982.

English:EPODD-7-2-75

[ET94] Paul M. English and Raman Tenneti. Interleaf active documents. Electronic Publishing —Origination, Dissemination, and Design, 7(2):75–87, June 1994. CODEN EPODEU. ISSN 0894-3982.

Furuta:EPODD-7-1-1

[FB94] Richard K. Furuta and David F.

Brailsford. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 7 (1):1–2, March 1994. CODEN EPODEU. ISSN 0894-3982.

Feiner:EPODD-3-1-29

[Fei90] Steven K. Feiner. Authoring large hypermedia documents with IGD. Electronic Publishing —Origination, Dissemination, and Design, 3(1):29–46, February 1990. CODEN EPODEU. ISSN 0894-3982.

Francois:EPODD-8-2/3-63

[FFE95] Patricia François, Philippe Futtersack, and Christophe Espert. SGML/HyTime repositories and object paradigms. Electronic Publishing—Origination, Dissemination, and Design, 8 (2/3):63–79, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Filgueiras:EPODD-6-4-507

[FL93] Miguel Filgueiras and José Paulo Leal. Representation and manipulation of music documents in ScEX. Electronic Publishing —Origination, Dissemination, and Design, 6(4):507, December 1993. CODEN EPODEU. ISSN 0894-3982.

Furuta:EPODD-2-4-211

[FPS89] Richard K. Furuta, Catherine Plaisant, and Ben Shneiderman. Automatically transforming regularly structured documents into Hypertext. Electronic Publishing—Origination, Dissemination, and Design, 2 (4):211–229, December 1989. CODEN EPODEU. ISSN 0894-3982.

Furuta:EPODD-1-1-19

[FQA88] R. Furuta, V. Quint, and J. André. Interactively editing structured documents. Electronic Publishing—Origination, Dissemination, and Design, 1 (1):19–44, April 1988. CODEN EPODEU. ISSN 0894-3982.

Furuta:EPODD-3-4-179

[FS90] Richard K. Furuta and P. David Stotts. A functional metastructure for hypertext models and systems. Electronic Publishing—Origination, Dissemination, and Design, 3(4):179– 205, November 1990. CODEN EPODEU. ISSN 0894-3982.

Furuta:EPODD-5-1-19

[Fur92] Richard K. Furuta. Important papers in the history of document preparation systems: basic sources. Electronic Publishing—Origination, Dissemination, and Design, 5(1):19– 44, March 1992. CODEN EPODEU. ISSN 0894-3982.

Feng:EPODD-6-4-361

[FW93] An Feng and Toshiro Wakayama. SIMON: A grammarbased transformation system for structured documents. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):361–372, December 1993. CODEN EPODEU. ISSN 0894-3982.

Fankhauser:EPODD-6-4-447

[FX93] Peter Fankhauser and Yi Xu.

MarkItUp! an incremental approach to document structure recognition. Electronic Publishing—Origination, Dissemination, and Design, 6(4):447– 456, December 1993. CODEN EPODEU. ISSN 0894-3982.

Groves:EPODD-6-4-315

[GB93] Michael J. Groves and David F.

Brailsford. Separate compilation of structured documents. Electronic Publishing—Origination, Dissemination, and Design, 6(4):315–326, December 1993. CODEN EPODEU. ISSN 0894-3982.

Gonczarowski:EPODD-2-3-157

[GP89] Jakob Gonczarowski and On G. Paradise. InScript—a C-like preprocessor for PostScript. Electronic Publishing—Origination, Dissemination, and Design, 2(3):157–167, October 1989. CODEN EPODEU. ISSN 0894-3982.

Hall:EPODD-7-3-113

[Hal94] Wendy Hall. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 7 (3):113–115, September 1994. CODEN EPODEU. ISSN 0894-3982.

Hammersley:EPODD-5-2-97

[Ham92] P. Hammersley. EP-odds and ends: A curriculum in electronic publishing. Electronic Publishing—Origination, Dissemination, and Design, 5(2):97– 102, June 1992. CODEN EPODEU. ISSN 0894-3982.

Hansen:EPODD-3-1-3

[Han90] Bo Stig Hansen. A functionbased formatting model. Electronic Publishing—Origination, Dissemination, and Design, 3 (1):3–28, February 1990. CO-DEN EPODEU. ISSN 0894-3982.

Haralambous:EPODD-6-3-145

[Har93] Yannis Haralambous. Parametrization of PostScript fonts through —an alternative to Adobe Multiple Master Fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):145– 157, September 1993. CODEN EPODEU. ISSN 0894-3982.

Haralambous:EPODD-7-4-197

[Har94a] Yannis Haralambous. Typesetting Khmer. Electronic Publishing—Origination, Dissemination, and Design, 7(4):197– 215, December 1994. CODEN EPODEU. ISSN 0894-3982.

Harrison:EPODD-7-3-117

[Har94b] Leon Harrison. A review of multimedia technology and dissemination systems. Electronic Publishing—Origination, Dissemination, and Design, 7(3):117– 146, September 1994. CODEN EPODEU. ISSN 0894-3982.

Haralambous:EPODD-8-2/3-111

[Har95] Yannis Haralambous. The traditional Arabic typecase extended to the Unicode set of glyphs. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 111–123, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Habusha:EPODD-3-2-65

[HB90] Uri Habusha and Daniel Berry.

Vi.iv, a bi-directional version of the vi full-screen editor. Electronic Publishing—Origination, Dissemination, and Design, 3 (2):65–91, May 1990. CODEN EPODEU. ISSN 0894-3982.

Heeman:EPODD-5-3-143

[Hee92] Frans C. Heeman. Granularity in structured documents. Electronic Publishing—Origination, Dissemination, and Design, 5 (3):143–155, September 1992. CODEN EPODEU. ISSN 0894-3982.

Hersch:EPODD-7-4-195

[Her94] Roger D. Hersch. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 7(4):195–196, December 1994. CODEN EPODEU. ISSN 0894-3982.

Hatzimanikatis:EPODD-7-1-35

[HGC94] A. Hatzimanikatis, I. Gaviotis, and D. Christodoulakis. Distributed Documents: an architecture for open distributed hypertext. Electronic Publishing —Origination, Dissemination, and Design, 7(1):35–48, March 1994. CODEN EPODEU. ISSN 0894-3982.

Herz:EPODD-6-3-261

[HH93] Jacky Herz and Roger D. Hersch. Analysing character shapes by string matching techniques. Electronic Publishing—Origination, Dissemination, and Design, 6(3):261–272, September 1993. CODEN EPODEU. ISSN 0894-3982.

Herz:EPODD-7-4-251

[HH94] J. Herz and R. D. Hersch. Towards a universal auto-hinting system for typographic shapes.

Electronic Publishing—Origination, Dissemination, and Design, 7(4):251–260, December 1994. CODEN EPODEU. ISSN 0894-3982.

Haake:EPODD-7-2-89

[HHR94] Anja Haake, Christoph Hüser, and Klaus Reichenberger. The Individualized Electronic Newspaper: an example of electronic publication. Electronic Publishing—Origination, Dissemination, and Design, 7(2):89– 111, June 1994. CODEN EPODEU. ISSN 0894-3982.

Hu:EPODD-6-4-457

[HI93] Tao Hu and Rolf Ingold. A mixed approach toward an efficient logical structure recognition from document images. Electronic Publishing—Origination, Dissemination, and Design, 6(4):457–468, December 1993. CODEN EPODEU. ISSN 0894-3982.

Harrison:EPODD-2-4-193

[HM89] Michael A. Harrison and Ethan V. Munson. On integrated bibliography processing. Electronic Publishing—Origination, Dissemination, and Design, 2(4):193–209, December 1989. CODEN EPODEU. ISSN 0894-3982.

Harrison:EPODD-4-1-43

[HM91] Michael A. Harrison and Ethan V. Munson. Numbering document components. Electronic Publishing—Origination, Dissemination, and Design, 4 (1):43–60, March 1991. CO-DEN EPODEU. ISSN 0894-3982.

Huser:EPODD-6-4-519

[HM93] Christoph Hüser and Wiebke Möhr. Colophon. Electronic Publishing—Origination, Dissemination, and Design, 6(4): 519–520, December 1993. CO-DEN EPODEU. ISSN 0894-3982.

Hussain:EPODD-6-3-171

[HP93] Fiaz Hussain and Michael L. V. Pitteway. Rasterizing the outlines of fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):171– 181, September 1993. CODEN EPODEU. ISSN 0894-3982. Hayashi:EPODD-6-4-423

[HS93] Koichi Hayashi and Akifumi Sekijima. Mediating interface between hypertext and structured documents. Electronic Publishing—Origination, Dissemination, and Design, 6(4): 423–434, December 1993. CO-DEN EPODEU. ISSN 0894-3982.

Itoh:EPODD-6-3-195

[IO93] Koichi Itoh and Yoshio Ohno. A curve fitting algorithm for character fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):195, September 1993. CODEN EPODEU. ISSN 0894-3982.

Jones:EPODD-3-4-207

[JS90] Tricia Jones and Ben Shneiderman. Examining usability for a training-oriented hypertext: Can hyper-activity be good? Electronic Publishing—Origination, Dissemination, and Design, 3(4):207– 225, November 1990. CODEN EPODEU. ISSN 0894-3982.

Kacmar:EPODD-8-1-31

[Kac95] Charles J. Kacmar. A process approach for providing hypermedia services to existing, nonhypermedia applications. Electronic Publishing—Origination, Dissemination, and Design, 8 (1):31–48, March 1995. CO-DEN EPODEU. ISSN 0894-3982.

Kaelbling:EPODD-3-2-93

[Kae90] Michael J. Kaelbling. On improving SGML. Electronic Publishing—Origination, Dissemination, and Design, 3(2):93–98, May 1990. CODEN EPODEU. ISSN 0894-3982.

Karow:EPODD-4-3-151

[Kar91] Peter Karow. Digital punch cutting. Electronic Publishing—Origination, Dissemination, and Design, 4(3):151– 170, September 1991. CODEN EPODEU. ISSN 0894-3982.

Kumar:EPODD-8-2/3-235

[KFA95] Vijay Kumar, Richard Furuta, and Robert B. Allen. Interactive interfaces for knowledgerich domains. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 235–246, June/September 1995.

CODEN EPODEU. ISSN 0894-3982.

King:EPODD-8-2/3-95

[Kin95] P. R. King. Modelling multimedia documents. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 95–110, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Klassen:EPODD-6-3-183

[Kla93] R. Victor Klassen. Variable width splines: a possible font representation? Electronic Publishing—Origination, Dissemination, and Design, 6(3):183– 194, September 1993. CODEN EPODEU. ISSN 0894-3982.

Knuth:EPODD-3-2-99

[Knu90] Donald E. Knuth. A note on digitized angles. Electronic Publishing—Origination, Dissemination, and Design, 3 (2):99–104, May 1990. CODEN EPODEU. ISSN 0894-3982.

Kokula:EPODD-7-4-217

[Kok94] Michael Kokula. Automatic generation of script font ligatures based on curve smoothness optimization. Electronic Publishing—Origination, Dissemination, and Design, 7(4): 217–229, December 1994. CO-DEN EPODEU. ISSN 0894-3982.

Kuikka:EPODD-6-4-373

[KP93] Eila Kuikka and Martti Penttonen. Transformation of structured documents with the use of grammar. Electronic Publishing—Origination, Dissemination, and Design, 6(4):373– 383, December 1993. CODEN EPODEU. ISSN 0894-3982.

Kuikka:EPODD-8-4-319

[KP95] E. Kuikka and M. Penttonen.

Transformation of structured documents. Electronic Publishing—Origination, Dissemination, and Design, 8(4):319– 341, December 1995. CODEN EPODEU. ISSN 0894-3982.

Kuikka:EPODD-8-2/3-181

[KS95] E. Kuikka and A. Salminen.

Filtering structured documents in the SYNDOC environment. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):181–193, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Lovegrove:EPODD-8-2/3-207

[LB95] William S. Lovegrove and David F. Brailsford. Document analysis of PDF files: Methods, results and implications. Electronic Publishing—Origination, Dissemination, and Design, 8 (2/3):207–220, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

DeLange:EPODD-6-3-241

[LEB93] Rudi W. De Lange, Hendry L.

Esterhuizen, and Derek Beatty. Performance differences between Times and Helvetica in a reading task. Electronic Publishing—Origination, Dissemination, and Design, 6(3):241– 248, September 1993. CODEN EPODEU. ISSN 0894-3982.

Levy:EPODD-6-4-339

[Lev93] David M. Levy. Document reuse and document systems. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):339–348, December 1993. CODEN EPODEU. ISSN 0894-3982.

Lamb:EPODD-6-1-23

[LL93] David Alex Lamb and Margaret Anne Lamb. Separation of concerns for indexing. Electronic Publishing—Origination, Dissemination, and Design, 6 (1):23–34, March 1993. CO-DEN EPODEU. ISSN 0894-3982.

Lucarella:EPODD-1-2-105

[Luc88] D. Lucarella. A search strategy for large document bases. Electronic Publishing—Origination, Dissemination, and Design, 1 (2):105–116, September 1988. CODEN EPODEU. ISSN 0894-3982.

Lucarella:EPODD-4-4-183

[Luc91] Dario Lucarella. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 4 (4):183–184, December 1991. CODEN EPODEU. ISSN 0894-3982.

Mandel:EPODD-6-1-3

[Man93] Ladislas Mandel. Developing an awareness of typographics letterforms. Electronic Publishing —Origination, Dissemination, and Design, 6(1):3–22, March 1993. CODEN EPODEU. ISSN 0894-3982.

Mamrak:EPODD-4-1-27

[MB91] Sandra A. Mamrak and J. A.

Barnes. Considerations for the preparation of SGML document type definitions. Electronic Publishing—Origination, Dissemination, and Design, 4 (1):27–42, March 1991. CO-DEN EPODEU. ISSN 0894-3982.

McQueen:EPODD-6-3-117

[MB93] Clyde D. McQueen III and Raymond G. Beausoleil. Infinifont: a parametric font generation system. Electronic Publishing—Origination, Dissemination, and Design, 6(3):117– 132, September 1993. CODEN EPODEU. ISSN 0894-3982.

Morris:EPODD-6-4-349

[MBM93] Robert A. Morris, Edward M.

Blachman, and Charles Meyer. A constraint-based editor for linguistic scholars. Electronic Publishing—Origination, Dissemination, and Design, 6(4): 349–360, December 1993. CO-DEN EPODEU. ISSN 0894-3982.

McCleery:EPODD-5-2-55

[McC92] Alistair McCleery. Teaching electronic publishing: a Scottish example. Electronic Publishing—Origination, Dissemination, and Design, 5(2):55–61, June 1992. CODEN EPODEU. ISSN 0894-3982.

Munson:EPODD-8-2/3-125

[Mun95] Ethan V. Munson. A new presentation language for structured documents. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 125–138, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Murata:EPODD-8-4-295

[Mur95]  Makoto Murata.  File for-

mat for documents containing both logical structures and layout structures. Electronic Publishing—Origination, Dissemination, and Design, 8(4):295– 317, December 1995. CODEN EPODEU. ISSN 0894-3982.

Nicholas:EPODD-6-2-93

[NR93] Charles K. Nicholas and Linda H. Rosenberg. Canto: a hypertext data model. Electronic Publishing—Origination, Dissemination, and Design, 6(2):93–113, June 1993. CODEN EPODEU. ISSN 0894-3982.

Nicholas:EPODD-5-3-105

[NW92] Charles K. Nicholas and Lawrence A. Welsch. On the interchangeability of SGML and ODA. Electronic Publishing—Origination, Dissemination, and Design, 5 (3):105–130, September 1992. CODEN EPODEU. ISSN 0894-3982.

Oakley:EPODD-1-2-79

[ON88] A. L. Oakley and A. C. Norris. Page description languages: Development, implementation and standardization. Electronic Publishing—Origination, Dissemination, and Design, 1(2): 79–96, September 1988. CO-DEN EPODEU. ISSN 0894-3982.

Pozzi:EPODD-5-3-131

[PCS92] Silvano Pozzi, Augusto Celentano, and Luisa Salemme. ALIVE: A distributed live-link documentation system. Electronic Publishing—Origination, Dissemination, and Design, 5 (3):131–142, September 1992. CODEN EPODEU. ISSN 0894-3982.

Pausch:EPODD-3-4-227

[PD90] R. Pausch and J. Detmer.

Node popularity as a hypertext browsing aid. Electronic Publishing—Origination, Dissemination, and Design, 3(4):227– 234, November 1990. CODEN EPODEU. ISSN 0894-3982.

Pickering:EPODD-5-1-49

[Pic92] Adrian Pickering. Book review:

David Collier, Collier’s Rules for Desktop Design and Typography. Electronic Publishing —Origination, Dissemination, and Design, 5(1):49–50, March 1992. CODEN EPODEU. ISSN 0894-3982.

Pickering:EPODD-7-1-51

[Pic94] Adrian Pickering. Book review: Antoni Diller, L^{A}TEX Line by Line, Tips and Techniques for Document Processing. Electronic Publishing—Origination, Dissemination, and Design, 7 (1):51–52, March 1994. CO-DEN EPODEU. ISSN 0894-3982.

Park:EPODD-6-2-67

[PM93] Seung Woon Park and Seung Ryoul Maeng. Structure extraction and automatic hinting of Chinese outline characters. Electronic Publishing —Origination, Dissemination, and Design, 6(2):67–91, June 1993. CODEN EPODEU. ISSN 0894-3982.

Pollard:EPODD-3-3-155

[Pol90] Richard Pollard. Hypertext presentation of thesauri used in online searching. Electronic Publishing—Origination, Dissemination, and Design, 3(3):155– 172, August 1990. CODEN EPODEU. ISSN 0894-3982.

Quint:EPODD-6-4-313

[QHM93] Vincent Quint, Christoph Hüser, and Wiebke Möhr. Editorial: Special issue: Proceedings of EP’94, the Electronic Publishing, Document Manipulation, and Typography Conference. Electronic Publishing—Origination, Dissemination, and Design, 6(4):313– 314, December 1993. CODEN EPODEU. ISSN 0894-3982.

Quint:EPODD-7-2-53

[Qui94] Vincent Quint. Editorial. Electronic Publishing—Origination, Dissemination, and Design, 7 (2):53–54, June 1994. CODEN EPODEU. ISSN 0894-3982.

Quint:EPODD-7-2-55

[QV94] Vincent Quint and Irène Vatton. Making structured documents active. Electronic Publishing—Origination, Dissemination, and Design, 7(2):55–74, June 1994. CODEN EPODEU. ISSN 0894-3982.

Richy:EPODD-8-2/3-81

[RA95] Hélène Richy and Jacques André. Typographic sheets and structured documents. Electronic Publishing—Origination, Dissemination, and Design, 8 (2/3):81–93, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Rada:EPODD-3-3-125

[Rad90] Roy Rada. Hypertext writing and document reuse: the role of a semantic net. Electronic Publishing—Origination, Dissemination, and Design, 3 (3):125–140, August 1990. CO-DEN EPODEU. ISSN 0894-3982.

Rahtz:EPODD-5-1-51

[Rah92a] Sebastian Rahtz. Book review:

Norbert Schwartz, Introduction to TEX. Electronic Publishing —Origination, Dissemination, and Design, 5(1):51–52, March 1992. CODEN EPODEU. ISSN 0894-3982.

Rahtz:EPODD-5-1-45

[Rah92b] Sebastian Rahtz. Book review:

Philip Crookes, Ventura Publisher. Electronic Publishing —Origination, Dissemination, and Design, 5(1):45–46, March 1992. CODEN EPODEU. ISSN 0894-3982.

Reid:EPODD-1-1-55

[Rei88] B. K. Reid. The USENET cookbook—an experiment in electronic publishing. Electronic Publishing—Origination, Dissemination, and Design, 1(1): 55–76, April 1988. CODEN EPODEU. ISSN 0894-3982.

Richy:EPODD-7-1-21

[Ric94] Hélène Richy. A hypertext electronic index based on the Grif structured editor. Electronic Publishing—Origination, Dissemination, and Design, 7 (1):21–34, March 1994. CO-DEN EPODEU. ISSN 0894-3982.

Rostek:EPODD-6-4-495

[RMF93] Lothar Rostek, Wiebke Möhr, and Dietrich H. Fischer. Weaving a web: the structure and creation of an object network representing an electronic reference work. Electronic Publishing—Origination, Dissemination, and Design, 6(4):495– 505, December 1993. CODEN EPODEU. ISSN 0894-3982.

Rizk:EPODD-7-3-179

[RML94] Antoine Rizk, Francis Malezieux, and Alain Leger. Using the MHEG standard in the hypermedia system Multicard. Electronic Publishing—Origination, Dissemination, and Design, 7 (3):179–193, September 1994.

CODEN EPODEU. ISSN 0894-3982.

Rossiter:EPODD-3-3-141

[RSH90] B. N. Rossiter, T. J. Sillitoe, and M. A. Heather. Database support for very large hypertexts. Electronic Publishing —Origination, Dissemination, and Design, 3(3):141–154, August 1990. CODEN EPODEU. ISSN 0894-3982.

Roisin:EPODD-6-4-327

[RV93] Cécile Roisin and Irène Vatton.

Merging logical and physical structures in documents. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):327–337, December 1993. CODEN EPODEU. ISSN 0894-3982.

Srouji:EPODD-5-4-163

[SB92] Johny Srouji and Daniel Berry. Arabic formatting with

ditroff/ffortid. Electronic

Publishing—Origination, Dissemination, and Design, 5(4): 163–208, December 1992. CO-DEN EPODEU. ISSN 0894-3982.

Scheermesser:EPODD-6-3-207

[SB93] Thomas Scheermesser and Olof Bryngdahl. Digital halftoning with texture control. Electronic Publishing—Origination, Dissemination, and Design, 6 (3):207–212, September 1993. CODEN EPODEU. ISSN 0894-

Smith:EPODD-8-2/3-153

[SB95] Philip N. Smith and David N. Brailsford. Towards structured, block-based PDF. Electronic Publishing—Origination, Dissemination, and Design, 8 (2/3):153–165, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Salton:EPODD-5-1-1

[SBA92] Gerard Salton, Chris Buckley, and James Allan. Automatic structuring of text files. Electronic Publishing—Origination, Dissemination, and Design, 5 (1):1–17, March 1992. CODEN EPODEU. ISSN 0894-3982.

Smith:EPODD-6-4-481

[SBE^{+}93] Philip N. Smith, David F. Brailsford, David R. Evans, Leon Harrison, Steve G. Probets, and Peter E. Sutton. Journal publishing with Acrobat: the CAJUN project. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):481–493, December 1993. CODEN EPODEU. ISSN 0894-3982.

Savoy:EPODD-4-2-87

[SD91] Jacques Savoy and Daniel Desbois. Information retrieval in hypertext systems: an approach using Bayesian networks. Electronic Publishing—Origination, Dissemination, and Design, 4 (2):87–108, June 1991. CODEN EPODEU. ISSN 0894-3982.

Sennhauser:EPODD-6-3-273

[Sen93] René Sennhauser. Improving the recognition accuracy of text recognition systems using typographical constraints. Electronic Publishing—Origination, Dissemination, and Design, 6 (3):273–282, September 1993. CODEN EPODEU. ISSN 0894-3982.

Stotts:EPODD-4-2-119

[SF91] P. David Stotts and Richard K.

Furuta. Hypertext 2000: Databases or documents. Electronic Publishing—Origination, Dissemination, and Design, 4 (2):119–121, June 1991. CO-DEN EPODEU. ISSN 0894-3982.

Sullow:EPODD-6-4-413

[SGBO^{+}93] Klaus Süllow, Ingrid Gabel-Becker, Marlies Ockenfeld, Wolfgang Putz, and Gisela Roth. MultiMedia Forum: an interactive online journal. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):413–422, December 1993. CODEN EPODEU. ISSN 0894-3982.

Schnase:EPODD-6-1-35

[SLH^{+}93] John L. Schnase, John J.

Leggett, David L. Hicks, Peter J. Nuernberg, and J. Alfredo Sánchez. Design and implementation of the HB1 hyperbase management system. Electronic Publishing—Origination, Dissemination, and Design, 6 (1):35–63, March 1993. CO-DEN EPODEU. ISSN 0894-3982.

Stamm:EPODD-6-3-219

[Sta93a] Beat Stamm. Dynamic regularisation of intelligent outline fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):219– 230, September 1993. CODEN EPODEU. ISSN 0894-3982.

Stamm:EPODD-6-3-159

[Sta93b] Beat Stamm. Object-orientation and extensibility in a fontscaler. Electronic Publishing—Origination, Dissemination, and Design, 6(3):159– 170, September 1993. CODEN EPODEU. ISSN 0894-3982.

Sanderson:EPODD-4-4-205

[SvR91] Mark Sanderson and C. J. van Rijsbergen. NRT: news retrieval tool. Electronic Publishing—Origination, Dissemination, and Design, 4(4):205– 217, December 1991. CODEN EPODEU. ISSN 0894-3982.

Terry:EPODD-3-2-105

[TB90] Douglas B. Terry and Donald G. Baker. Active Tioga documents: an exploration of two paradigms. Electronic Publishing—Origination, Dissemination, and Design, 3(2):105– 122, May 1990. CODEN EPODEU. ISSN 0894-3982.

vanOssenbruggen:EPODD-8-2/3-51

[vES95] Jacco van Ossenbruggen, Anton Eliëns, and Bastian Schönhage. Web applications and SGML.

Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):51–62, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Weir:EPODD-5-1-47

[Wei92] Thomas E. Weir, Jr. Book review: A. Rizk, N. Streitz, and J. André, editors, First European Conference on Hypertext. Electronic Publishing—Origination, Dissemination, and Design, 5 (1):47–48, March 1992. CO-DEN EPODEU. ISSN 0894-3982.

Wonneberger:EPODD-7-1-3

[Won94] Reinhard Wonneberger. TEX in an industrial environment. Electronic Publishing—Origination, Dissemination, and Design, 7(1):3–19, March 1994. CODEN EPODEU. ISSN 0894-3982.

Warmer:EPODD-2-2-65

[WV89] J. Warmer and S. Van Egmond.

The implementation of the Amsterdam SGML parser. Electronic Publishing—Origination, Dissemination, and Design, 2 (2):65–90, July 1989. CODEN EPODEU. ISSN 0894-3982.

Warmer:EPODD-4-1-3

[WvV91] Jos Warmer and Hans van Vliet. Processing SGML documents. Electronic Publishing —Origination, Dissemination, and Design, 4(1):3–26, March 1991. CODEN EPODEU. ISSN 0894-3982.

Wang:EPODD-8-2/3-167

[WW95] Xinxin Wang and Derick Wood. Xtable — A tabular editor and formatter. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 167–179, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Zapf:EPODD-6-3-283

[Zap93] Hermann Zapf. About microtypography and the hz-program. Electronic Publishing—Origination, Dissemination, and Design, 6(3):283–288, September 1993. CODEN EPODEU. ISSN 0894-3982.

Zapf:EPODD-7-4-261

[Zap94] Herman Zapf. Call for foundation of a ‘Sir Francis Drake Society’. Electronic Publishing—Origination, Dissemination, and Design, 7(4):261– 263, December 1994. CODEN EPODEU. ISSN 0894-3982.

Zeggel:EPODD-6-3-213

[ZB93] Thomas Zeggel and Olof Bryngdahl. Error diffusion on an adaptive raster. Electronic Publishing—Origination, Dissemination, and Design, 6(3):213– 218, September 1993. CODEN EPODEU. ISSN 0894-3982. Zhang:EPODD-8-2/3-221

[ZD95] Qin Zhang and John M. Danskin. A pattern-based lossy compression scheme for document images. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):

221–233, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Zramdini:EPODD-6-3-249

[ZI93] Abdel Wahab Zramdini and Rolf Ingold. Optical font recognition from projection profiles. Electronic Publishing—Origination, Dissemination, and Design, 6(3):249–260, September 1993. CODEN EPODEU. ISSN 0894-3982.